OpenStudy (anonymous):

What is the value of the expression? 6! / 3! x 3! I know "x" isn't the proper thing to do for multiplication but I can't do the other sign. A. 1/3 B. 1 C.20 D. 80 @dhiraj123

OpenStudy (anonymous):

is it 6i/3i*3i

OpenStudy (anonymous):

6!/3!*3!

OpenStudy (anonymous):

what does this sign mean(!)

OpenStudy (anonymous):

I don't know, actually. That's just how the problem is.

OpenStudy (anonymous):

omg marry me! :DD

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i am sorry eflatt13 i dont know either

OpenStudy (math&ing001):

The sign (!) stands for 'factorial'. The factorial of a number is the product of all positive integers less than or equal to it. Example : 5! = 1x2x3x4x5

OpenStudy (math&ing001):

In your case : \[\frac{ 6! }{ 3! * 3! }=\frac{ 1*2*3*4*5*6 }{ (1*2*3)*(1*2*3) }=\frac{ 4*5*6 }{ 6 }=20\]

OpenStudy (anonymous):

isnt denominator going to be 36

OpenStudy (math&ing001):

Yes, and that's the common factor between the nominator and the denominator. You get 36*20/36=20.